OpenStudy (anonymous):

Please help me simplify this: (sin^4x -cos^4x)/(sin^2x -cos^2x)

OpenStudy (anonymous):

\[\frac{ \sin ^{4x}-\cos ^{4x} }{ \sin ^{2x}-\cos ^{2x} }\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait no

terenzreignz (terenzreignz):

\[\Large \frac{ \sin ^{4}x-\cos ^{4}x }{ \sin ^{2}x-\cos ^{2}x }\]

OpenStudy (anonymous):

yes that is what it is teren

terenzreignz (terenzreignz):

There's something to love about the difference of two squares... or two quartics, as long as the exponents are even.

OpenStudy (anonymous):

so use difference of 2 squares on the top and bottom?

terenzreignz (terenzreignz):

Yeah, but, pro-tip, don't mind the bottom. Focus on the top, factor away...

OpenStudy (anonymous):

alright thank you very much

terenzreignz (terenzreignz):

You left me hanging :P Do me a favour... when you're done, post your answer. I'm curious :)

OpenStudy (anonymous):

will do, and keep an eye out for the ton of identiy problems i'm stuck on

OpenStudy (anonymous):

i got 1 as my answer

terenzreignz (terenzreignz):

Very good.

OpenStudy (anonymous):

thanks a lot! :)

terenzreignz (terenzreignz):

No problem. Heck, you did most of the work yourself ^.^